what is the % of the S enantiomer in a sample of carvone that has a specific rotatin of -20, given that the specific rotation of R-carvone is -61?
Answer #1

Optical purity % = (specific rotation of sample/specific rotation of pure sample) x 100

As we have sample rotation -20, and R-carvone rotation -61, we can say that R-carvone is in excess in the sample.

optical purity % = (-20/-61) x 100 = 32.79 %

The sample has 32.79% excess of R-carvone

Remaining = 100 - 32.79 = 67.21 %

% of S-enantiomer in sample would be = 67.21/2 = 33.60 %